# DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD

### v1

@@ B.4 Proof of Theorem 3
Assume that throughout $t$ iterations where $\tau\ll t\ll\mathbf{T}$, the constants $a$, $b$, $\delta$, and $\tau$ remain fixed. Then the average iteration time $T_{avg}:=\frac{TC_{t}}{t}$ satisfies the following approximation:
$T_{avg}\approx\max{\left\{\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1},\frac{\delta S_{g}}{a},T_{\text{comp}}\right\}},$ (18)
where the approximation error is upper bounded by $O(\frac{1}{t})$.
@@
$\left\{\begin{matrix}TC_{k}=TM_{k}+b\\ TS_{k+1}=T_{\text{comp}}+max\left\{TC_{k-\tau},TS_{k}\right\}\\ TM_{k+1}=\frac{\delta S_{g}}{a}+max\left\{TM_{k},TS_{k+1}\right\}\end{matrix}\right.\quad,k=0,1,2...\quad.$ (19)
@@
$TC_{t}=TM_{t}+b=\frac{\delta S_{g}}{a}+tT_{\text{comp}}+b,$
@@
$\left|TC_{t}-tT_{\text{comp}}\right|=\frac{\delta S_{g}}{a}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
@@
$\left|TC_{t}-t\cdot\frac{\delta S_{g}}{a}\right|=T_{\text{comp}}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
@@
{$TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=\frac{\delta S_{g}}{a}+iT_{\text{comp}},\end{aligned}\right.\quad\text{for }i=1,\dots,\tau+1;$ (20)
{$TS_{i}&=\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\\ TM_{i}&=2\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\end{aligned}\right.\quad\text{for }i=\tau+2,\dots,2\tau+2.$
@@
$A_{k}$ :=$:=\left\{(TM_{k(\tau+1)+d},TS_{k(\tau+1)+d})\mid d=1,\dots,\tau+1\right\},$
$A_{k}+a$ :=$:=\left\{(TM_{k(\tau+1)+d}+a,TS_{k(\tau+1)+d}+a)\mid d=1,\dots,\tau+1\right\},\quad\forall a\in\mathbb{R}.$
@@
Let $t=m(\tau+1)+d$ for $d=1,\dots,\tau+1$, then the cumulative time $TC_{t}$ satisfies:
$TC_{t}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+dT_{\text{comp}}+\frac{\delta S_{g}}{a}+b,$
$\left|TC_{t}-t\left(\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)\right|$ $=d\left(T_{\text{comp}}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+\frac{\delta S_{g}}{a}+b$
@@
This confirms that $TC_{t}$ deviates from its average rate by at most a bounded constant, further supporting the regularity and predictability of the update pattern in this intermediate-delay regime.
@@
{$TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=i\frac{\delta S_{g}}{a}+T_{\text{comp}},\end{aligned}\right.\quad$ $i=1,\dots,\tau+1;$
{$TS_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-2)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-1)\tfrac{\delta S_{g}}{a},\end{aligned}\right.\quad$ $i=\tau+2,\dots,2\tau+2;$
{$TS_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-3)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-2)\tfrac{\delta S_{g}}{a},\end{aligned}\right.\quad$ $i=2\tau+3,\dots,3\tau+3.$
@@
Now, let $t=m(\tau+1)+d$ with $d\in\{1,\dots,\tau+1\}$. The cumulative computation time $TC_{t}$ satisfies:
$TC_{t}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+d\cdot\frac{\delta S_{g}}{a}+T_{\text{comp}}+b,$
$\left|TC_{t}-t\cdot\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right|$ $=d\left(\frac{\delta S_{g}}{a}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+T_{\text{comp}}+b$
@@
$T^{\prime}_{\text{avg}}$ =$=\begin{cases}\dfrac{T_{\text{comp}}+\frac{\delta S_{g}}{a}+b}{\tau+1},&\tau<\tau_{\text{thres}},\\ \max\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\},&\tau\geq\tau_{\text{thres}},\end{cases}\quad\text{with}\quad\left|TC_{t}-tT^{\prime}_{\text{avg}}\right|\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\},$
@@
$\left|T_{\text{avg}}-T^{\prime}_{\text{avg}}\right|\leq\frac{b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}}{t}=O\left(\frac{1}{t}\right).$
@@ Appendix C Addendum to Evaluation Experiments
C.1 Experimental Environment
The experiments are conducted on an Ubuntu $24.04$ LTS server with an Intel Xeon Gold $6230$ CPU and $3$ Nvidia A40404040 GPUs, each with $48$GB memory. The Python version is $3.10.16$, and other used libraries are all based on the Python version. We use PyTorch $2.5.1$ with CUDA $12.4$ as the ML toolkit. We use gloo as the communication backend.
C.2 Detailed Experimental Settings
Experiment tasks: The experimental setup encompasses three task types spanning computer vision (CV) and natural language processing (NLP), utilizing both traditional (CNN [29]) and state-of-the-art transformer-based models (ViT [30] and GPT [31]). The CNN architecture includes two convolutional layers followed by two fully connected layers, used widely in previous works [9, 20]. The ViT base model features approximately 86 million parameters, while the smallest GPT variant contains 124 million parameters.
To demonstrate the superiority of our design, we evaluate four representative tasks: FashionMNIST [32] and CIFAR-$10$ with CNN (denoted as CNN@FMNIST and CNN@CIFAR-$10$), ImageNet [33] using ViT (denoted as ViT@ImageNet) and Wikitext [34] on GPT (denoted as GPT@Wikitext). For CNN@FMNIST and CNN@CIFAR-$10$, the learning rate is $0.01$, and the batch size is $32$. For ViT@ImageNet and GPT@Wikitext, the learning rates are $0.01$ and $0.1$, respectively, with batch sizes of $40$ and $5$.
Baselines: We compare D-SGD, Accordion, DGA, and vanilla DD-EF-SGD as baselines. D-SGD is the case $\tau=0,\delta=1$, and While it does not suffer from accuracy degradation, it converges the slowest in high-latency and low-bandwidth scenarios when measured in terms of end-to-end time. Accordion [35] is the SOTA adaptive gradient compression algorithm, which sets small $\delta$ in the critical iterations and sets large $\delta$ if not [9]. The compressor we use in all baselines is Top-$k$, which transmits the largest $k$ elements of gradients. The reason why we select Top-$k$ is its strong empirical performance, widespread adoption in DML, and straightforward implementation. DGA is a state-of-the-art algorithm originally designed for Federated Learning [18], which first introduces delay staleness to reduce the negative effect of the high latency on Federated Learning. We set $K=1$ in DGA to close the infrequent communication. CocktailSGD [21] uses the fixed $\tau$ and $\delta$. The hyperparameters are chosen by DeCo-SGD with $E=\infty$.
Network environment: We simulated challenging network environments (static high latency $\geq$ $100$ms, dynamic low bandwidth, average bandwidth $\leq$ $1$Gbps), where the dynamic volatility of the network can be referred to the following figure, which gives the network environment in the given environment (Section 5.2), and we discuss the superiority of the DeCo-SGD algorithm in different network conditions (Section 5.4).
C.3 Network Conditions of Section 5.2
As shown in Fig. LABEL:fig:network, the network conditions exhibit dynamic bandwidth fluctuations across all tasks (CNN@FMNIST, CNN@CIFAR-10, ViT@ImageNet, and GPT@Wikitext), and the compression ratio $\delta$ is adaptively adjusted over time under a fixed 200 ms transmission delay with $\tau=2$.
C.4 Detailed Curves of Section 5.3 to Verify the Scalability of DeCo-SGD
The detail of Fig. 5 in the main text of the paper corresponds to Fig. 9 in the Appendix.
@@ C.2 Detailed Experimental Settings
C.5 Detailed Results of Section 5.4
Fig. 10 in the Appendix provides a visual counterpart to the data presented in Table 1 in the main text. Table 3 extends more details in Table 1
Table 3: Comparison of training time (s) under different methods and network conditions. $a$ is the average bandwidth and $b$ is the end-to-end latency. $\tau^{*}$ and $\delta^{*}$ are caculated by DeCo-SGD. The numbers in parentheses represent the speedup compared to DeCo-SGD.
Model @DataSet a(Gbps), b(s) $\tau^{*}$, $\delta^{*}$ D-SGD Accordion DGA CocktailSGD DeCo-SGD
GPT @Wikitext $0.1,0.1$ $2,0.02$ $6396.95$ (4.90×4.90$\times$4.90 ×) $1955.28$ (1.50×1.50$\times$1.50 ×) $5506.78$ (4.22×4.$2$) $1525.21$ (1.17×1.17$\times$1.17 ×) $\mathbf{1306.29}$
$0.5,0.1$ $2,0.10$ $1571.37$ (2.44×2.$4$) $829.06$ (1.29×1.29$\times$1.29 ×) $1346.34$ (2.09×2.09$\times$2.09 ×) $712.93$ ($1\times1$) $\mathbf{642.79}$
$0.1,1.0$ $3,0.02$ $7232.88$ (5.02×5.02$\times$5.02 ×) $2510.26$ (1.74×1.74$\times$1.74 ×) $5764.60$ (4.00×4.$0$) $1683.11$ (1.17×1.17$\times$1.17 ×) $\mathbf{1440.59}$
$0.5,1.0$ $3,0.10$ $2363.73$ (2.37×2.37$\times$2.37 ×) $1325.12$ (1.33×1.$3$) $1562.93$ (1.57×1.57$\times$1.57 ×) $1132.63$ (1.14×1.14$\times$1.14 ×) $\mathbf{997.25}$
ViT @ImageNet $0.1,0.1$ $2,0.03$ $1373.84$ (4.85×4.85$\times$4.85 ×) $440.14$ (1.55×1.$5$) $1183.68$ (4.18×4.18$\times$4.18 ×) $364.77$ (1.29×1.29$\times$1.29 ×) $\mathbf{283.06}$
$0.5,0.1$ $2,0.14$ $446.20$ (2.64×2.64$\times$2.64 ×) $274.60$ (1.62×1.62$\times$1.62 ×) $385.60$ (2.28×2.28$\times$2.28 ×) $219.39$ (1.30×1.30$\times$1.30 ×) $\mathbf{169.13}$
$0.1,1.0$ $3,0.03$ $1530.57$ (3.85×3.85$\times$3.85 ×) $604.67$ (1.52×1.52$\times$1.52 ×) $1287.03$ (3.24×3.24$\times$3.24 ×) $505.91$ (1.27×1.27$\times$1.27 ×) $\mathbf{397.16}$
$0.5,1.0$ $3,0.14$ $631.15$ (2.67×2.67$\times$2.67 ×) $373.79$ (1.58×1.58$\times$1.58 ×) $444.38$ (1.88×1.$8$) $278.57$ (1.18×1.18$\times$1.18 ×) $\mathbf{235.95}$
(a) ViT@ImageNet bandwidth=$100$Mbps latency=$100$ms
(b) ViT@ImageNet bandwidth=$100$Mbps latency=$1000$ms
(c) ViT@ImageNet bandwidth=$500$Mbps latency=$100$ms
(d) ViT@ImageNet bandwidth=$500$Mbps latency=$1000$ms
(e) GPT@Wikitext bandwidth=$100$Mbps latency=$100$ms
(f) GPT@Wikitext bandwidth=$100$Mbps latency=$1000$ms
(g) GPT@Wikitext bandwidth=$500$Mbps latency=$100$ms
(h) GPT@Wikitext bandwidth=$500$Mbps latency=$1000$ms
@@
Appendix D Others
D.1 UsedResources and Licenses
All external assets used in this paper are properly credited, and their licenses and terms of use have been explicitly acknowledged and respected. Specifically, we make use of several open-source resources as follows:
  1. 1.
    ViT model, released under the MIT License.
  2. 2.
    GPT2 model from Hugging Face, released under the MIT License.
  3. 3.
    docker-tc, used to simulate network conditions, released under the MIT License.
All of these assets were used strictly for research purposes in accordance with their respective licenses. We have cited the corresponding authors or repositories in our references where appropriate and ensured full compliance with their terms of use.
D.2 Limitations and Boarder Impact
Limitations: (1) Although DeCo-SGD currently supports only the D-SGD and its variants, we believe that D-SGD is applicable to a wide range of tasks, including NLP, CV and Objective Detection. If D-SGD is properly parameterized, its performance is no worse than popular optimizers in LLM training such as Adam [52]. (2) Device heterogeneity (different bandwidth/latency per node) is not considered, as current theoretical frameworks (based on convex optimization) cannot analyze such scenarios without prohibitive parameter search costs. Nevertheless, our approach already achieves $80\%$ throughput gains over SOTA optimizer [21] in dynamic networks, leaving limited room for further improvement via heterogeneity optimization.
Boarder Impact: Our work establishes theoretical foundations for LLM pre-training in low-bandwidth, high-latency networks. While employing lossy training techniques, we provide theoretical guarantees of training efficiency. This could significantly reduce the infrastructure costs for LLM pre-training, promoting broader participation in AI research and development.

### v2

@@ B.5 Proof of Theorem 3
Under Assumptions from 5 to 8, the average iteration time $T_{avg}$ satisfies the following approximation:
where the approximation error is upper bounded by $O(\frac{1}{E})$, where $E$ is the analysis window.
@@
$\left\{\begin{matrix}TC_{k}=TM_{k}+b\\ TS_{k+1}=T_{\text{comp}}+max\left\{TC_{k-\tau},TS_{k}\right\}\\ TM_{k+1}=\frac{\delta S_{g}}{a}+max\left\{TM_{k},TS_{k+1}\right\}\end{matrix}\right.\quad,k=0,1,2.$$\quad.$ (24)
@@
$TC_{E}=TM_{E}+b=\frac{\delta S_{g}}{a}+ET_{\text{comp}}+b,$
@@
$\left|TC_{E}-ET_{\text{comp}}\right|=\frac{\delta S_{g}}{a}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
@@
$\left|TC_{E}-E\cdot\frac{\delta S_{g}}{a}\right|=T_{\text{comp}}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
@@
$\left\{\begin{aligned} TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=\frac{\delta S_{g}}{a}+iT_{\text{comp}},\end{aligned}\right.\quad\text{for }i=1,\dots,\tau+1;$ (25)
$\left\{\begin{aligned} TS_{i}&=\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\\ TM_{i}&=2\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\end{aligned}\right.\quad\text{for }i=\tau+2,\dots,2\tau+2.$
@@
$A_{k}$ $=\left\{(TM_{k(\tau+1)+d},TS_{k(\tau+1)+d})\mid d=1,\dots,\tau+1\right\},$
$A_{k}+a$ $=\left\{(TM_{k(\tau+1)+d}+a,TS_{k(\tau+1)+d}+a)\mid d=1,\dots,\tau+1\right\},\quad\forall a\in\mathbb{R}.$
@@
Let $E=m(\tau+1)+d$ for $d=1,\dots,\tau+1$, then the cumulative time $TC_{E}$ satisfies:
$TC_{E}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+dT_{\text{comp}}+\frac{\delta S_{g}}{a}+b,$
$\left|TC_{E}-E\left(\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)\right|$ $=d\left(T_{\text{comp}}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+\frac{\delta S_{g}}{a}+b$
@@
This confirms that $TC_{E}$ deviates from its average rate by at most a bounded constant, further supporting the regularity and predictability of the update pattern in this intermediate-delay regime.
@@
$\left\{\begin{aligned} TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=i\frac{\delta S_{g}}{a}+T_{\text{comp}},\end{aligned}\right$$\quad$ $i=1,\dots,\tau+1;$
$\left\{\begin{aligned} TS_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-2)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-1)\tfrac{\delta S_{g}}{a},\end{aligned}\right$$\quad$ $i=\tau+2,\dots,2\tau+2;$
$\left\{\begin{aligned} TS_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-3)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-2)\tfrac{\delta S_{g}}{a},\end{aligned}\right$$\quad$ $i=2\tau+3,\dots,3\tau+3.$
@@
Now, let $E=m(\tau+1)+d$ with $d\in\{1,\dots,\tau+1\}$. The cumulative computation time $TC_{E}$ satisfies:
$TC_{E}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+d\cdot\frac{\delta S_{g}}{a}+T_{\text{comp}}+b,$
$\left|TC_{E}-E\cdot\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right|$ $=d\left(\frac{\delta S_{g}}{a}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+T_{\text{comp}}+b$
@@
[…]
@@
$\left|T_{\text{avg}}-T^{\prime}_{\text{avg}}\right|\leq\frac{b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}}{E}=O\left(\frac{1}{E}\right).$
@@ Appendix C Addendum to Evaluation Experiments
C.1 Compression ratio and stelaness curves in Section 5.3
As shown in Fig. 9, the network conditions exhibit dynamic bandwidth fluctuations across all tasks (CNN@FMNIST, CNN@CIFAR-10, ViT@ImageNet, and GPT@Wikitext), and the compression ratio $\delta$ is adaptively adjusted over time under a fixed 200 ms transmission delay with $\tau=2$.
(a) ResNet@CIFAR-10
(b) ViT@ImageNet
(c) GPT@Wikitext
Figure 9: $\delta$ varying over time for tasks: ResNet@CIFAR-10, ViT@ImageNet, and GPT@Wikitext, with $\tau=2$.
C.2 Detailed Curves of Section 5.4 to Verify the Scalability of DeCo-SGD
Fig. 10 provide the detailed accuracy/ppl curves in Fig. 8,
@@
C.3 Detailed Results of Section 5.5
Fig. 11 provides a visual counterpart to the data presented in Table 3 in the main text.
(a) ViT@ImageNet bandwidth=100100Mbps latency=100100ms
(b) ViT@ImageNet bandwidth=100100Mbps latency=10001000ms
(c) ViT@ImageNet bandwidth=500500Mbps latency=100100ms
(d) ViT@ImageNet bandwidth=500500Mbps latency=10001000ms
(e) GPT@Wikitext bandwidth=100100Mbps latency=100100ms
(f) GPT@Wikitext bandwidth=100100Mbps latency=10001000ms
(g) GPT@Wikitext bandwidth=500500Mbps latency=100100ms
(h) GPT@Wikitext bandwidth=500500Mbps latency=10001000ms
@@
C.4 Experiment Results in non-IID settings
The non-IID partition strategy is distribution-based label imbalance, which allocates a proportion of samples of each label according to Dirichlet distribution, with the concentration parameter set to 0.5. The results demonstrate that under non-IID settings, DeCo-SGD consistently outperforms other methods.
C.5 Experiment Results with other sparsifiers
We compare DeCo-SGD with other optimizers under another sparsifier Random-K, which randomly selects k% gradient elements to transmit. The results below show that DeCo-SGD consistently outperforms other methods with other sparsifiers.
Table 5: Comparison of training time for DeCo-SGD and other optimizers under the Random-K sparsifier, which randomly selects k% gradient elements to transmit.
Model
@DataSet
a(Gbps),
b(s)
D-SGD DAGC Accordion DC2 DGA CocktailSGD DeCo-SGD
GPT
@Wikitext
$0.1,0.1$ $7276.34\pm 49.33$ ($5.13\times$) $2283.53\pm 12.55$ ($1.61\times$) $2246.34\pm 12.24$ ($1.59\times$) $2244.73\pm 12.21$ ($1.58\times$) $6108.13\pm 35.23$ ($4.31\times$) $1732.25\pm 12.58$ ($1.22\times$) $\mathbf{1417.55\pm 11.23}$
$0.5,0.1$ $1785.64\pm 12.55$ ($2.41\times$) $969.11\pm 7.12$ ($1.30\times$) $921.27\pm 7.56$ ($1.24\times$) $917.54\pm 7.42$ ($1.23\times$) $1515.61\pm 12.48$ ($2.04\times$) $815.52\pm 6.45$ ($1.10\times$) $\mathbf{741.35\pm 5.73}$
$0.1,1.0$ $8179.46\pm 54.32$ ($4.76\times$) $2841.85\pm 20.07$ ($1.65\times$) $2812.31\pm 20.35$ ($1.63\times$) $2827.36\pm 20.91$ ($1.64\times$) $6517.52\pm 34.66$ ($3.79\times$) $2018.44\pm 14.65$ ($1.17\times$) $\mathbf{1718.41\pm 13.46}$
$0.5,1.0$ $2736.48\pm 19.64$ ($2.31\times$) $1576.29\pm 12.98$ ($1.32\times$) $1566.14\pm 12.11$ ($1.32\times$) $1570.48\pm 12.07$ ($1.32\times$) $1775.52\pm 13.29$ ($1.50\times$) $1459.16\pm 11.63$ ($1.23\times$) $\mathbf{1186.33\pm 9.27}$
ViT
@ImageNet
$0.1,0.1$ $1273.93\pm 9.32$ ($4.64\times$) $567.73\pm 3.73$ ($2.06\times$) $545.18\pm 3.45$ ($1.98\times$) $537.85\pm 3.14$ ($1.95\times$) $1324.88\pm 10.26$ ($4.82\times$) $368.39\pm 1.82$ ($1.34\times$) $\mathbf{274.76\pm 1.45}$
$0.5,0.1$ $434.16\pm 2.15$ ($2.34\times$) $364.22\pm 1.80$ ($1.93\times$) $347.22\pm 1.73$ ($1.87\times$) $353.70\pm 1.79$ ($1.91\times$) $467.97\pm 2.10$ ($2.52\times$) $224.09\pm 1.38$ ($1.21\times$) $\mathbf{185.90\pm 0.68}$
$0.1,1.0$ $1724.74\pm 10.47$ ($4.22\times$) $672.44\pm 3.21$ ($1.64\times$) $654.28\pm 3.09$ ($1.60\times$) $644.02\pm 3.14$ ($1.57\times$) $1367.39\pm 8.25$ ($3.34\times$) $555.26\pm 3.11$ ($1.36\times$) $\mathbf{409.18\pm 2.48}$
$0.5,1.0$ $677.23\pm 4.21$ ($2.78\times$) $489.41\pm 2.21$ ($2.00\times$) $467.39\pm 2.07$ ($1.92\times$) $478.93\pm 1.85$ ($1.96\times$) $599.93\pm 2.83$ ($2.46\times$) $283.89\pm 1.62$ ($1.17\times$) $\mathbf{243.52\pm 1.02}$